### v3

# One Shot Learning as Instruction Data Prospector for Large Language Models

Abstract
Aligning large language models (LLMs) with human is a critical step in effectively utilizing their pre-trained capabilities across a wide array of language tasks. Current instruction tuning practices often rely on expanding dataset size without a clear strategy for ensuring data quality, which can inadvertently introduce noise and degrade model performance. To address this challenge, we introduce Nuggets, a novel and efficient methodology that employs one shot learning to select high-quality instruction data from expansive datasets. Nuggets assesses the potential of individual instruction examples to act as effective one shot examples, thereby identifying those that can significantly enhance diverse task performance. Nuggets utilizes a scoring system based on the impact of candidate examples on the perplexity of a diverse anchor set, facilitating the selection of the most beneficial data for instruction tuning. Through rigorous testing on two benchmarks, including MT-Bench and Alpaca-Eval, we demonstrate that instruction tuning with the top 1% of Nuggets-curated examples substantially outperforms conventional methods that use the full dataset. These findings advocate for a data selection paradigm that prioritizes quality, offering a more efficient pathway to align LLMs with humans.
00footnotetext: https://github.com/pldlgb/nuggets
1 Introduction
Large language models (LLMs) [5, 32, 3, 15, 4, 50, 22] have demonstrated impressive capabilities [48, 37, 29] across a wide range of language task through the scaling of model size and data. Based on the pre-training, it is necessary to further align them to better follow human instructions. The process involves supervised fine-tuning on input-output pairs, known as instruction tuning [31]. Instruction tuning is an indispensable step, that not only activates the useful knowledge gained by LLMs during pre-training, but also enables them to interact with humans in a natural conversational manner.
Considerable efforts in instruction tuning have focused on collecting larger [47, 10, 36, 46], diversity [36, 39, 19, 44], and complex [53, 49] datasets, typically through human crowd-sourcing [33, 9, 30, 1, 40] or stealing from larger models [45, 41, 8, 53, 49, 61]. Although there is an increase in the size of datasets used for instruction tuning, some works [63, 7, 6, 24] indicate smaller but valuable datasets tend to be more effective in harnessing the capability of LLMs. Blindly expanding the quantity of instruction data without ensuring quality might introduce noise and lead to hallucination issues [59, 60]. However, there are no standard criteria for selecting instruction data [25, 16, 51, 56]. As shown in Figure 1, the common practice is to rely on empirical methods for selection [52], which is biased, to determine data combination, adjusting based on end results. This trial-and-error approach increases the alignment costs for models. We believe optimal instruction combinations exist within the vast data available, but an efficient, low-cost method to identify them is less explored.
To contribute to the understanding of this topic, we introduce Nuggets, a simple and efficient method that leverages LLMs themselves as data explorers by one shot (in-context) learning, selecting the beneficial data from vast instructional datasets. Intuitively, an instruction example is worth training if it promotes that task as a few shot example of a specific task. If it can facilitate many tasks, it will be worth being treated as a prime data focus. Another potential argument comes from the fact that some work has shown that in-context learning [11, 2, 55, 18, 21, 43] utilizes prompting to implicitly fine-tune the model, while instruction tuning is via gradient descent, and it is promising that we can leverage the performance of in-context learning to predict the effects of instruction tuning. Concretely, we first select a set covering multiple tasks as the anchor set and the remaining examples as the candidate set. One example from the candidate set is sequentially selected to serve as a one shot example of in-context learning, and then scored by observing its impact on the perplexity of each anchor example. This score can infer the dependency between anchor and candidate examples and serve as a reference standard for data selection.
Figure 1: The comparison between our Nuggets and previous empirical methods. Unlike empirical methods (blue area), Nuggets method (orange area) can directly sample a gold subset to provide for model fine-tuning.
To validate the efficacy of the proposed Nuggets, we conduct extensive evaluations on two well-known benchmarks, MT-Bench [62] and Alpaca-Eval [26], we choose the most popular and powerful language model, LLaMA [42], as our base model. Empirical results demonstrate that, after employing the Nuggets’ data filtering strategy, there is a significant improvement compared to the vanilla fine-tuning. Notably, fine-tuning the LLM with just the top 1% of highest-scoring examples yielded better results than using the entire instruction dataset. This encourages a focus on the quality and strategic composition of data rather than sheer volume.
Before delving into details, we clearly emphasize our contribution as follows.
  • In this paper, we introduce Nuggets, a method that adaptively assesses the quality of instruction examples using large language models. Nuggets enables us to extract the most valuable data from a vast pool of instruction data for the purpose of fine-tuning.
  • With the proposed Nuggets, a series of superior model fine-tuning results are achieved by using only 1% instruction examples. This provides strong evidence for the effectiveness of our strategy compared to using entire 100% examples.
  • Comprehensive experiments indicate that our assumptions regarding golden instructions are well-founded. This insight may potentially provide some guidance for future endeavors in data quality screening.
2 Related Work
Instruction Tuning
Recent works have introduced a series of techniques that aim at refining large language models (LLMs), showcasing their ability to generalize effectively to instructions not encountered before. For instance, T5 [35] pioneered the initial effort of training various natural language processing (NLP) tasks in a unified text-to-text format. FLAN [47] introduced the novel concept of instruction tuning, aiming to improve zero shot task performance by transforming NLP tasks into natural language instructions during model training. Furthermore, InstructGPT [33] handled a wide array of human-created instructions encompassing diverse forms and a broad range of task types, tailored for real-world user scenarios. As OpenAI did not release the source code for these remarkable projects, Alpaca [41, 34] and Vicuna [8] took the initiative to delve into open-domain instruction tuning using the open-source LLM LLaMA [42]. In particular, Alpaca employed a dataset of 52k instructions generated from a small seed set of manually written instructions, comprising only 175 examples. Meanwhile, Vicuna utilized 70k user-shared conversations with ChatGPT, gathered from “ShareGPT.com”.
Instruction Construction
The fine-tuning instruction datasets by previous methods are often created manually or tailored to specific tasks. To alleviate the issue of extensive human annotations and manual data gathering, various semi-automated techniques have emerged. Self-Instruct [45] selected a handful of instances randomly from the initial task pool and used them as demonstrations to instruct a language model in generating new instructions along with corresponding input-output pairs. Evol-Instruct [53] progressively modified the original instruction in a step-by-step manner, allowing for precise control over the difficulty and complexity of the generated instructions. Tree-Instruct [61], in contrast to Self-Instruct or Evol-Instruct, guided LLMs by instructing them to append a specified number of new nodes to the semantic tree of an existing instruction rather than directly manipulating the text sequence. On the other hand, some research focuses on enhancing the performance of language models by utilizing fewer but higher-quality instruction examples. LIMA [63] demonstrated remarkably strong performance by strategically selecting a thousand high-quality data points for learning. InstructMining [6] introduced a set of selected natural language indicators for evaluating instruction-following data quality. However, this method requires the data to be divided into multiple bins and the models to be fully trained, which cannot assess the quality of individual examples at a fine-grained level. Moreover, ALPAGASUS [7] utilized the capabilities of an external and powerful model, ChatGPT, to directly evaluate each example. Although this method has been proven effective, it does not account for the inherent differences in each model that is being fine-tuned. Instead, it predominantly relies on the preferences of the ChatGPT model. Although [24] proposed a self-guided method for selecting data for instruction tuning, it still requires preliminary fine-tuning of the model, introducing uncertainty into subsequent operations.
Figure 2: The illustration of the framework of our Nuggets. Note that we do not directly let the model generate answers for assessment. Instead, we calculate the model’s logit scores on the ground truth answers as zero shot scores or one shot scores.
3 Nuggets
Motivation
As shown in Figure 1, historically, the refinement of instructional data for fine-tuning has predominantly been guided by empirical methodologies. This encompasses the application of heuristic rules, the scrutiny of human experts, and the iterative adaptation of data based on feedback gleaned from the model’s performance. The training of the trial-and-error method consumes a significant amount of human and computational resources. In this paper, we propose that Large Language Models (LLMs) can implicitly function as circumspect data explorers through in-context learning, which can further guide themselves to identify the necessary training samples. Based on this capability, we introduce Nuggets, a method that can efficiently select beneficial data from vast instructional datasets.
Overview
The framework of our Nuggets is illustrated in Figure 2. Firstly, we evaluate the proficiency of large language models on a variety of tasks using a predefined set of tasks, denoted as the zero shot score. Then, we take each example from the instruction dataset as an one shot prompt, concatenating it in front of the predefined tasks, and recalculate the model’s completion level for these tasks, referred to as the one shot score. By leveraging the difference between one shot and zero shot scores, we can compute the golden score for each instruction. Once we have the golden scores for all instructions, we can select the highest-scoring subset as the golden subset, which is then directly provided to the model for fine-tuning.
3.1 Algorithm Details
Zero Shot Score
[…]
$s^{j}_{\text{zsl}}=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|{T}_{j},{w}_{1}^{\text{A}_{j}},{w}_{2}^{\text{A}_{j}},\ldots,{w}_{i-1}^{\text{A}_{j}};\textsf{LLM}),$ (1)
where $L$ is the number of words of the ground-truth answer A. The score ${s}^{j}_{\text{zsl}}$ is employed to signify the extent of the model’s proficiency on the $j$-th task. A higher ${s}^{j}_{\text{zsl}}$ denotes superior model performance on the $j$-th task, whereas a lower ${s}^{j}_{\text{zsl}}$ implies inferior performance. Therefore, we can acquire the model’s performance across $m$ tasks, denoted as:
$\bm{S}_{\text{zsl}}=[{s}^{1}_{\text{zsl}},{s}^{2}_{\text{zsl}},\ldots,{s}^{m-1}_{\text{zsl}},{s}^{m}_{\text{zsl}}].$ (2)
One Shot Score
With an instruction tuning dataset $\mathcal{D}$, we aim to identify a set of examples $\mathcal{D}_{\text{gold}}$ that most closely align with the golden instructions. For each example $\mathbf{z}_{k}=[\text{Instruction}^{\text{Q}}_{k}~{}(\text{IQ}_{k}),\text{Instruction}^{\text{A}}_{k}~{}(\text{IA}_{k})]$, we initially perform implicit instruction tuning on the base model using that specific example. Subsequently, we employ the meta-optimized updated model to conduct another round of testing on the tasks within the task pool. That is,
${s}^{j}_{\text{iit}}(\mathbf{z}_{k})$ $=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|\underbrace{\text{IQ}_{k},\text{IA}_{k}}_{\textit{One Shot Prompt}},{T}_{j},{w}^{\text{A}_{j}}_{1},{w}^{\text{A}_{j}}_{2},\ldots,{w}^{\text{A}_{j}}_{i-1};\texttt{LLM}),$ (3)
where $\text{IQ}_{k}$ and $\text{IA}_{k}$ can be considered analogous to $\mathbf{X}_{\text{ins}}$ in Eq. (6), while T and ${w}^{\text{A}}_{1},{w}^{\text{A}}_{2},\ldots,{w}^{\text{A}}_{i-1}$ can be regarded as $\mathbf{X}_{\text{test}}$. Similarly, we can obtain the performance of the model after meta-optimization across $m$ different tasks:
$\bm{S}^{k}_{\text{iit}}=[{s}^{1}_{\text{iit}}(\mathbf{z}_{k}),{s}^{2}_{\text{iit}}(\mathbf{z}_{k}),\ldots,{s}^{m-1}_{\text{iit}}(\mathbf{z}_{k}),{s}^{m}_{\text{iit}}(\mathbf{z}_{k})].$ (4)
Afterward, we use the Golden Score (GS) to reflect the impact of this instruction tuning example on the base model. The GS of the example $\mathbf{z}_{k}$ is calculated as
$\text{GS}(\mathbf{z}_{k})=\frac{1}{m}\sum_{i=1}^{m}\mathbbm{I}\left[{s}^{i}_{\text{iit}}(\mathbf{z}_{k})>{s}^{i}_{\text{zsl}}\right]\ \in\ [0,1],$ (5)
where $\mathbbm{I}[\cdot]$ is the indicator function. At a high level, the GS measures the increment of performance improvement of the model after one shot learning through the given instruction.
In this work, we calculate the GS for each instruction example, enabling us to generate a ranking of the scores for the entire set of examples. Our objective is to explicitly fine-tune the base model using only a small subset of the most critical examples, specifically focusing on the examples with high golden scores, to achieve better results than using the whole dataset.
Figure 3: The distribution of the golden score for the Alpaca instruction dataset. In a total of 52,002 cases, there are 9,549 instructions with a gold score less than 0.5, indicating that these data have a side effect on overall task completion. Besides, there are 7,524 instructions with a gold score greater than 0.8, suggesting that the model improves the task completion rate through one shot learning from these data, which can be considered as high-quality instruction data.
4 Experiments
4.1 Experimental Setup
Instruction Dataset
We exploit the Alpaca dataset [41]. It is an important resource in the open-source community for instruction tuning, which is constructed by employing the self-instruct [45] method to distill instruction data from text-davinci-003. The success of this dataset in fine-tuning the LLaMA model has sparked a series of explorations into instruction fine-tuning [23, 20, 54, 28, 64, 12].
Predefined Task Set
The predefined task set under consideration is associated with the earlier cited task pool, which is used to compute golden scores for instructions. These data primarily serve to assess the model’s capacity to generalize across a variety of tasks, allowing for considerable leeway in data selection. The predefined task set is deemed adequate provided it includes a sufficient quantity of data and encapsulates a wide array of tasks. Given that the Alpaca dataset naturally displays these characteristics, we have randomly selected 1,000 examples from it to form the predefined task set.
Evaluation Datasets
Numerous endeavors are dedicated to evaluating the abilities of LLMs [27, 14, 13, 17, 26, 62]. However, it remains an ongoing and unresolved issue. Presently, a commonly adopted strategy involves employing the robust language model GPT-4 as the evaluator [26, 62, 8, 7]. In this framework, we employ two methods to assess the model’s capabilities. The first involves rating the responses generated by the model on a scale of 1 to 10. A representative dataset for this purpose is MT-Bench [62], which evaluates instruction-following ability across eight categories: writing, roleplay, extraction, reasoning, math, coding, STEM, and humanities. The second method involves comparing the model’s generated responses to Davinci-003 responses, using a well-known dataset called Alpaca-Eval [26]. This dataset uses “win_rate” as the evaluation metric.
Implementation Details
We choose the LLaMa-7b model as the foundational model for our experiments. To ensure a fair comparison, we also set the maximum input length for the models fine-tuned with the Alpaca dataset to be consistent with LLaMa, which is 2048. In the model fine-tuning phase, we employ the Adam optimizer with a learning rate of $2\times 10^{-5}$ and use a batch size of 64, and conduct training over three epochs. As for the model evaluation phase, we retained all the parameter settings from the original work [26, 62].
4.2 Experimental Results
The Alpaca dataset encompasses a total of 52,002 instruction examples, with the distribution of their golden scores illustrated in Figure 3. Out of these, 42,460 examples have a golden score exceeding 0.5. We have also chosen some examples that are closely aligned with the golden instructions, notably those with golden scores above 0.8 and 0.85. There are 7,525 examples with golden scores surpassing 0.8 and 619 with golden scores exceeding 0.85. The latter constitutes a mere 1% of the entire dataset.
We conduct instruction tuning on LLaMA using different subsets of examples with golden scores: less than 0.5, greater than 0.5, greater than 0.8, greater than 0.85, and the full dataset. The fine-tuned models are named $\text{Alpaca}_{\leq 0.5}$, $\text{Alpaca}_{>0.5}$, $\text{Alpaca}_{>0.8}$, $\text{Alpaca}_{>0.85}$, and $\text{Alpaca}_{\text{full}}$ respectively.
Table 1: Experimental results on Alpaca-Eval Benchmark.
Model Nums Helpful_Base Koala Self-instruct Oasst Vicuna Length Overall
LLaMA - 0.00 1.28 1.19 0.53 1.25 2,980 0.87
[…]
Alpaca${}_{>0.5}$ 42,460 24.03 20.51 28.57 29.78 15.00 413 26.06
Alpaca${}_{>0.8}$ 7,525 34.10 30.76 30.95 35.10 30.00 519 32.48
Alpaca${}_{>0.85}$ 619 37.20 26.90 25.00 29.30 22.50 617 28.20
Table 2: Experimental results on MT-Bench Benchmark.
Model Writing Roleplay Reasoning Math Coding Extraction STEM Humanities Overall
LLaMA 4.6 4.5 5.2 1.0 1.20 2.2 5.0 4.1 3.47
[…]
Alpaca${}_{>0.5}$ 8.3 5.7 3.5 1.1 1.7 5.0 6.6 7 4.86
Alpaca${}_{>0.8}$ 8.3 5.9 5.6 1.8 2.5 4.0 7.3 7.4 5.34
Alpaca${}_{>0.85}$ 6.6 6.3 4.9 1.0 2.3 3.3 6.3 7.3 4.87
Main Results
The experimental results are presented in Table 1 and Table 2. As expected, $\text{Alpaca}_{>0.8}$ produces the most impressive outcomes. This can be attributed to its ability to maintain an optimal balance between the volume and quality of the instructions it utilizes, leading to the most desirable results. We also note that incorporating lower-quality instructions adversely affected model fine-tuning. This trend is clear when we see that $\text{Alpaca}_{\leq 0.5}$ lagged behind $\text{Alpaca}_{\text{full}}$ in performance, while $\text{Alpaca}_{>0.5}$ shows a slight edge over $\text{Alpaca}_{\text{full}}$. Remarkably, $\text{Alpaca}_{>0.85}$, using only 1% of the dataset for fine-tuning, achieved results comparable to or even surpassing those of $\text{Alpaca}_{\text{full}}$. This underscores the efficacy of our data selection method.
Figure 4: The distribution of the golden score for the instruction dataset across different predefined task sets, along with the corresponding fine-tuning results on the Alpaca-Eval benchmark.
Ablation on Predefined Task Sets
To evaluate how different predefined task sets affect the selection of instruction data for fine-tuning, we include two additional predefined task set variations. One is randomly exampled from the Alpaca dataset but with a smaller task set size, which is limited to 100 examples. The other one entails clustering the Alpaca dataset into 100 clusters using the KMeans algorithm and selecting the centroids of each cluster as examples of the task set.
We use the two predefined sets to calculate golden scores for the Alpaca dataset separately. The distribution of golden scores is depicted in Figure 4. We select instruction data with golden scores less than or equal to 0.5, greater than 0.5, greater than 0.6, greater than 0.7, greater than 0.8, greater than 0.85, and greater than 0.9 for model fine-tuning, respectively. The results in Table 3 suggest that with random sampling, increasing the size of the task set can enhance the identification of high-quality instruction data. The logic behind this is that a larger encompasses a broader diversity of data, facilitating a more nuanced assessment of an instruction’s effect on model task generalization. However, a shift occurs when KMeans is employed to cherry-pick more distinct examples for the task set. With as few as 100 examples, KMeans outshines the results from 1,000 examples acquired through random sampling. In this instance, $\text{Alpaca}_{>0.8}$ delivered a superior performance with just 5,419 examples, compared to the 7,524 examples seen with $\text{Random}_{1000}$. This outcome also indirectly confirms the validity of our hypothesis regarding the definition of golden instructions.
Table 3: Experimental results on Alpaca-Eval Benchmark across different predefined task sets
[…]
Ablation on Instruction Set
To delve deeper into the generalization capabilities of Nuggets across varied directive datasets, we undertake a series of experiments utilizing the Alpaca-GPT4 dataset [34]. It generates instructional data from the powerful GPT-4 model [32], which is considered to have superior data quality. Additionally, it shares the same questions in instructions with the Alpaca dataset, which facilitates our direct comparison between the two.
Inspired by Table 3, we employ the KMeans algorithm on the Alpaca-GPT4 dataset to sample 100 examples, forming the predefined task set. Subsequently, we apply the Nuggets method to score all instructions in the Alpaca-GPT4 dataset with the golden score, as depicted in Figure 5. Compared to the Alpaca dataset, the Alpaca-GPT4 dataset boasts a significantly higher number of instructions with golden scores: 25,100 instructions exceed a score of 0.8, 16,943 surpass 0.85, and 4,250 instructions exceed 0.9. These numbers far exceed the corresponding high-scoring instructions in the Alpaca dataset. This also demonstrates that the golden score can serve as an absolute metric to assess the quality of instructional data. The results from model fine-tuning indicate that on the Alpaca-GPT4 dataset, conclusions align with those of previous experiments. The large language models fine-tuned on subsets with golden scores less than or equal to 0.5 exhibit the poorest performance, with a win rate of only 19.23% in the Alpaca-Eval benchmark. In contrast, the models fine-tuned on subsets with golden scores greater than 0.85 demonstrate superior performance, boasting a high win rate of 72.05%. This success can be attributed to the dual assurance of quantity and quality in this particular subset. It is worth emphasizing that fine-tuning on a small and high-quality dataset consistently and significantly outperforms the results of fine-tuning on the full dataset. Overall, the models fine-tuned using Alpaca-GPT4 significantly outperform those fine-tuned with Alpaca. This implicitly corroborates the superior quality of the Alpaca-GPT4 dataset compared to the Alpaca dataset.
Figure 5: The distribution of the golden score for the instruction dataset across different instruction sets, along with the corresponding fine-tuning results on the Alpaca-Eval benchmark. Both predefined task sets utilize KMeans to sample 100 examples from their respective instruction datasets.
Case Study
To evaluate Nuggets qualitatively, we also select some instruction examples for case studies, as shown in Figure 6. We observe that noisy instructions without outputs display extremely low golden score values. Such examples consistently score very low throughout the instruction dataset. Additionally, instructions with hallucinatory outputs and some that are a bit confusing also typically exhibit lower golden scores. In contrast, instructions with higher scores often possess fluent language and meticulous logic, and are oriented towards assisting humans in problem-solving. Therefore, we believe that the high golden scores can represent high-quality instruction tuning data.
Figure 6: Examples of instructions and their corresponding golden scores.
5 Discussion: One Shot Learning as Implicit Instruction Tuning
A widely acknowledged philosophy is that instruction tuning significantly enhances the task generalization of pre-trained models to perform various specific tasks [31, 57, 58, 38]. Therefore, we introduce the hypothesis of a golden instruction: the evaluation of an instruction’s effectiveness is determined by its impact on the model’s task generalization capability following the fine-tuning process. As the degree of improvement becomes more pronounced, the instruction approaches a classification as “golden instruction”.
Based on this hypothesis, a straightforward method is to fine-tune an independent model using an instruction example and then compare the performance of the fine-tuned model with the base model on a predefined dataset containing multiple tasks, to determine whether this example qualifies as a golden instruction. However, this way will result in an impractical number of fine-tuned models, equivalent to the number of instructions. Additionally, fine-tuning with only one example may introduce unstable updates to the model’s gradients, making it challenging to determine whether the example has genuinely been learned. Motivated by the inherent duality between attention and gradient descent [11, 2, 55, 18], we incorporate the meta-optimize mechanism into Transformer attention, replacing the need for actually fine-tuning the model.
[…]
$\approx\mathbf{W}_{V}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]\left(\mathbf{W}_{K}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]\right)^{\top}\textit{{Q}}$
$=\underbrace{\mathbf{W}_{V}\mathbf{X}_{\text{test}}(\mathbf{W}_{K}\mathbf{X}_{\text{test}})^{\top}}_{\textit{ Only test input.}}\textit{{Q}}+\underbrace{\mathbf{W}_{V}\mathbf{X}_{\text{ins}}(\mathbf{W}_{K}\mathbf{X}_{\text{ins}})^{\top}}_{\textit{Only instruction sample.}}\textit{{Q}}$
$=\mathbf{W}_{\text{zsl}}\ \textit{{Q}}+\Delta\mathbf{W}_{\text{iit}}\ \textit{{Q}}$
$=(\mathbf{W}_{\text{zsl}}+\Delta\mathbf{W}_{\text{iit}})\ \textit{{Q}},$ (6)
where $\sqrt{d_{\text{in}}}$ serves as a scaling factor. The term $\mathbf{W}_{V}\mathbf{X}_{\text{test}}(\mathbf{W}_{K}\mathbf{X}_{\text{test}})^{\top}$ could be denoted as $\mathbf{W}_{\text{zsl}}$, which represents the zero shot learning scenario where no instruction tuning is performed, since it solely focuses on the test input. In addition, the term $\mathbf{W}_{V}\mathbf{X}_{\text{ins}}(\mathbf{W}_{K}\mathbf{X}_{\text{ins}})^{\top}$ can be seen as implicit instruction tuning $\Delta\mathbf{W}_{\text{iit}}$ achieved via the meta-gradient [11, 55] derived from the instruction sample. Readers can refer to previous papers [11, 2, 18] for more details on implicit instruction tuning.
6 Conclusion
In this paper, we introduce Nuggets, a method that utilizes large language models to uncover more critical data for instruction tuning. It is based on one shot learning to detect the value of examples, which allows for efficient data selection and does not rely on additional annotation and cost. Benefiting from Nuggets, we find that even smaller subsets yield better instruction followings. Furthermore, we argue that our method emphasizes the need for data selection and provides insights for future instruction fine-tuning processes, encouraging fine-grained data combinations rather than blindly expanding quantities.
Future Work
While its efficacy has been confirmed in our experiments, there remain avenues for improvement. First, we can investigate including a diverse and compact task set in the golden scoring phase. The investigation will facilitate a more efficient model evaluation of instructional data and better identification of high-quality instructions for model fine-tuning. Second, due to resource constraints, the experiments of this work are limited to the LLaMA-7b model. While it is one of the most influential works in the large language model open-source community, further validation across a wider range of models is essential. Lastly, it is important to validate the proposed method on a more extensive set of instructional datasets.